# Jailbreaking Quantum Computers

### v1

@@ 4.3 Compound Attacks
The channel attacks and pulse attacks can be used as building blocks for compound attacks. In compound attacks, the individual channel attacks or pulse attacks are used as a building block to achieve higher-level attack goal. We list several interesting compound attacks that can be implemented by leveraging our attacks:
@@
  • Multi-Tenant Attack: entanglement can be also created in the multi-tenant environment. Though multi-tenant quantum computers are not supported yet, this paradigm is actively researched [35, 36, 37, 38], and our attacks can be applied in terms of multi-tenant use cases, especially with the channel attacks.
@@ 5 Attack Demonstration
5.1 Qubit Flip Attack Example
Given the extensive control over quantum circuits afforded by the attacks, one generic approach involves the attacker adding an extra layer of custom gate on all qubits either before measurement to manipulate final results or after the beginning to control input states. Here, we present one straightforward implementation for each attack to realize qubit flip. The implementations are shown in Figure 3, but the attack clearly extends beyond mere qubit flip.
5.1.1 Qubit Flip Attack using Channel Attacks
@@
Figure 3: Implementations of qubit flip with channel attacks and pulse attacks. The grey block in the channel attack section represents the custom gate that is manipulated in the attacks.
5.1.2 Qubit Flip Attack using Pulse Attacks
@@
(a) Quantum circuit of quantum teleportation. Alice (blue block) wants to transport her quantum state to Bob (green block), with the coupling gate provided by Eve (red block). Eve can tamper the pulses to achieve many goals, some of which are listed in the grey block.
@@ 6.2 Pulse Syntax Verification
We propose that to associate the quantum circuit with both the pulse configuration and the calibration data at the time the pulse configuration is calibrated, or identically the date on which the pulse configuration is calibrated. This associated data should be stored at some trusted authorities. Before users execute the quantum circuit or before the circuit publisher updates the pulse configuration, the input pulse configuration will be compared with the trusted pulse configuration, together with the analysis of the current calibration data from the quantum computer vendors or user custom experiments with the trusted calibration data.

### v2

@@ 4.3 Compound Attacks
The channel attacks and pulse attacks can be used as building blocks for compound attacks. In compound attacks, individual channel attacks or pulse attacks are used as a building block to achieve higher-level attack goals. We list several interesting compound attacks that can be implemented by leveraging our attacks:
@@
  • Multi-Tenant Attack: entanglement can also be created in the multi-tenant environment. Though multi-tenant quantum computers are not supported yet, this paradigm is actively researched [39, 40, 41, 42], and our attacks can be applied in terms of multi-tenant use cases, especially with the channel attacks.
@@ 5 Attack Demonstration
5.1 Toy Attack Example: Qubit Flip Attack
Given the extensive control over quantum circuits afforded by the attacks, one generic approach involves the attacker utilizing an extra layer of custom gates on all qubits either in the end to manipulate final results or at the beginning to control input states. This requires the modification of gate-level circuits, which requires attackers to be more powerful and is beyond the main threat model of this paper. This example serves to show one most straightforward way to perform the attack. Attackers may claim that the custom gate layer is for better fidelity, such as optimized readout pulses for lower measurement error or higher fidelity state preparation. Notice that level 1 victims have the lowest verification ability, and attackers can directly substitute the whole program. Here the goal is to present one simplest use case of the attack, and victims can check anything except the custom gate layer.
We present some straightforward implementations of the qubit flip attack. The implementations are shown in Figure 3, but the attack clearly extends beyond mere qubit flip.
5.1.1 Level 1 Victim: Qubit Flip Attack using Channel Attacks
@@
Figure 3: Implementations of qubit flip with channel attacks and pulse attacks. The gray block in the channel attack section represents the custom gate that is manipulated in the attacks.
5.1.2 Level 1 Victim: Qubit Flip Attack using Pulse Attacks
@@
(a) Quantum circuit of quantum teleportation. Alice (blue block) wants to transport her quantum state to Bob (green block), with the coupling gate provided by Eve (red block). Eve can tamper the pulses to achieve many goals, some of which are listed in the gray block.
@@ 6.2 Pulse Syntax Verification
We propose to associate the quantum circuit with both the pulse configuration and the calibration data at the time the pulse configuration is calibrated, or identically the date on which the pulse configuration is calibrated. This associated data should be stored at some trusted authorities. Before users execute the quantum circuit or before the circuit publisher updates the pulse configuration, the input pulse configuration will be compared with the trusted pulse configuration, together with the analysis of the current calibration data from the quantum computer vendors or user custom experiments with the trusted calibration data.
@@ 6.3 Pulse Semantics Verification
6.4 Evaluation
The defense framework was tested against all the attacks described in Section 5, and successfully detected all of them. However, the current implementation of the defense operates using an exact matching approach, allowing for variations in pulse amplitude within a specified tolerance range, as mentioned earlier. This design results in low precision but high recall for detecting malicious modifications: it effectively detects circuits where pulses or even gates have been tampered with, but it struggles to handle circuits that vary but generate identical results as the original circuits. For instance, a pulse intended for an X gate could be split into two pulses for an SX gate, and the current implementation would incorrectly flag this as a malicious circuit.
The trade-off between precision and recall can be adjusted a bit by modifying the tolerance range — higher tolerance increases precision while lowering recall. Usually, as circuits only need to be calibrated by simply adjusting the amplitude and the frequency to follow the current conditions of the environment, this exact matching approach proves useful. Nevertheless, a more advanced implementation that can accommodate variations in other pulse specifications is worth exploring and remains an area for future research.
@@ 8 Conclusion
Acknowledgments
The authors would like to thank IBM and Yale University for providing access to IBM’s superconducting devices. This work was also supported in part by NSF grants 2312754 and 2245344. The authors would like to thank the shepherd and the anonymous reviewers for their valuable feedback and comments used to improve the final version of this paper.
Appendix A Meta-Review
The following meta-review was prepared by the program committee for the 2025 IEEE Symposium on Security and Privacy (S&P) as part of the review process as detailed in the call for papers.
A.1 Summary
This paper explores implementation-level issues in the emerging world of quantum computing. Modern quantum computers implement quantum gates as microwave pulses, the exact specification of which vary according to calibration of different machines and over time. The authors show that, by precisely corrupting these pulse width specifications, attackers can make certain malicious changes to quantum circuits, without modifying the gate-level specification of the quantum circuit.
A.2 Scientific Contributions
  • Independent confirmation of important results with limited prior research.
  • Identifies an impactful vulnerability.
  • Establishes a new research direction.
A.3 Reasons for Acceptance
  1. 1.
    The paper is the first to propose this set of attacks, making it an interesting link in the chain stretching back to Ken Thompson’s 1984 observations on Trusting Trust.
  2. 2.
    The authors validated the effectiveness of attacks both via simulation and with devices, making it clear that the attacks are feasible.
  3. 3.
    The paper includes a good description of critical background that many readers may lack, despite the complexity of the subject matter.